OpenStudy (anonymous):

Solve the following system of equations. 9x + 4y = -30 3x + y = -9

OpenStudy (anonymous):

Solve for y y=-3x-9 substitute 9x+4(-3x-9)=-30 solve for y...

OpenStudy (anonymous):

Solve for x sorry...

OpenStudy (anonymous):

9x-12x-36=-30 -3x=6 x=-2

OpenStudy (anonymous):

Now substitute the x to solve for y

OpenStudy (anonymous):

Heyo